# A bound on the quantum value of all compiled nonlocal games

### v1

@@ 2.2. Algebras and representations
for all $A\in\mathscr{A}$. Moreover, $\lvert\nu_{\phi}\rangle$ is cyclic (meaning $\overline{\pi_{\phi}(\mathscr{A})\lvert\nu_{\phi}\rangle}=\mathcal{H}$) and thereby uniquely determined. We call $(\mathcal{H}_{\phi},\pi_{\phi},\lvert\nu_{\phi}\rangle)$ a GNS triple associated with $\phi$. For more information on $C^{*}$-algebras, we refer the reader to [Bla06].
@@
2.3. Classical and quantum computing
@@
For quantum computations, we will use the quantum circuit model. Here, computations correspond to the application of quantum circuits, which are unitary operators that operate on the Hilbert space $\mathcal{H}=(\mathbb{C}^{2})^{\otimes k}$ of some number $k$ of qubits and are given by the composition of unitary gates that each act nontrivially only on (for definiteness) one or two qubits (taken from some fixed universal gate set). The size of a quantum circuit is the number of gates used in the computation (we assume all qubits are acted upon by at least one gate). The qubits are typically split into input qubits and auxiliary qubits, which are assumed to be initialized in the $\lvert 0\rangle$ state, unless stated otherwise. If a classical outcome is desired, a subset of the qubits is measured after the unitary circuit has been applied. A quantum polynomial-time (QPT) algorithm consists of family of quantum circuits $\{C_{\lambda}\}_{\lambda\in\mathbb{N}}$ and a deterministic polynomial-time Turing machine that on input $1^{\lambda}$ outputs a description of $C_{\lambda}$. We can often interpret $\lambda$ as a problem size or as a security parameter.
@@
3. Nonlocal games and strategies
@@ 3.1. Nonlocal games
A (two-player) nonlocal game is a tuple $\mathcal{G}=(\mathcal{I}_{A},\mathcal{I}_{B},\mathcal{O}_{A},\mathcal{O}_{B},\mu,V)$ describing a scenario consisting of non-communicating players, Alice and Bob, interacting with a referee. In the game, the referee samples a pair of questions $(x,y)\in\mathcal{I}_{A}\times\mathcal{I}_{B}$ according to $\mu$, sending question $x$ to Alice and $y$ to Bob. Then, Alice (resp. Bob) returns answers $a$ (resp. $b$) to the referee, who computes the rule function $V$ on the question-answer pairs $(a,b,x,y)$ to determine if $V(a,b|x,y)=1$ they win, or $V(a,b|x,y)=0$ they lose. [^1] We use the notation $V(a,b|x,y)$ instead of $V(a,b,x,y)$ to emphasize that this represents the value of answers $a,b$ given questions $x,y$.
All the information about the game $\mathcal{G}$ is available to the player before the game. This allows them to decide on a strategy beforehand. However, once the game begins the players are not allowed to communicate. To the referee, the behavior of the players can be modeled by the probabilities $p(a,b|x,y)$ of answers $a,b$ given questions $x,y$ as determined by the strategy. The collection of numbers $\{p(a,b|x,y)\}_{a\in\mathcal{O}_{A},b\in\mathcal{O}_{B},x\in\mathcal{I}_{A},y\in\mathcal{I}_{B}}\in\mathbb{R}^{\mathcal{O}_{A}\times\mathcal{O}_{B}\times\mathcal{I}_{A}\times\mathcal{I}_{B}}$ is called a (bipartite) correlation. Thus, the probability of winning the game $\mathcal{G}$ under a strategy $S$, with correlations $p$, is given by
(3.1) $\omega(\mathcal{G},S)=\omega(\mathcal{G},p)=\sum_{x\in\mathcal{I}_{A},y\in\mathcal{I}_{B}}\sum_{a\in\mathcal{O}_{A},b\in\mathcal{O}_{B}}\mu(x,y)V(a,b|x,y)p(a,b|x,y).$
@@ 3.2. Strategies and correlations
Quantum strategies of particular interest are the entangled strategies. This is because if the state in the quantum strategy is unentangled, then the resulting correlation is always classical. It is easy to see that $C_{c}\subseteq C_{q}$, and the inclusion is in general strict, as follows from the existence of nontrivial Bell inequalities [CHSH69, Mer90, Per90].
The restriction to finite-dimensional quantum systems is not the only model. If one allows the Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ to be infinite-dimensional, one gets the set of spatial quantum correlations $C_{qs}$. Clearly, $C_{q}\subseteq C_{qs}$, and both sets are convex subsets $\mathbb{R}^{\mathcal{O}_{A}\times\mathcal{O}_{B}\times\mathcal{I}_{A}\times\mathcal{I}_{B}}$. It turns out that in general, the inclusion is strict [CS20] and neither set is closed [Slo17, DPP19, Col20, MR20, Bei21]. Moreover, both sets have the same closure, denoted by $C_{qa}$ and named the set of quantum approximable correlations.
@@
A (quantum) commuting operator strategy [^2] These are also called quantum commuting strategies in part of the literature. for a nonlocal game $\mathcal{G}$ consists of
@@
The set of commuting operator correlations $C_{qc}\subseteq\mathbb{R}^{\mathcal{O}_{A}\times\mathcal{O}_{B}\times\mathcal{I}_{A}\times\mathcal{I}_{B}}$ is always a closed convex subset [Fri12]. Since every quantum strategy is a commuting operator strategy by properties of the tensor product, it follows that $C_{qa}\subseteq C_{qc}$. When the Hilbert space $\mathcal{H}$ is finite-dimensional then any commuting operator strategy can also be seen as an ordinary quantum strategy, but in general, this is not so. In fact there exist commuting operator correlations which have no realization as a quantum strategy, and hence $C_{q}\subsetneq C_{qc}$ [Slo16]. Whether the correlation sets $C_{qa}$ and $C_{qc}$ were the same became known as Tsirelson’s Problem. This was recently resolved in the celebrated work $\mathsf{MIP}^{*}=\mathsf{RE}$ [JNV+21] by the construction of a correlation in $C_{qc}$ with no realization in $C_{qa}$. In turn, this implied a negative resolution to Connes’ Embedding Problem, following [Oza13, JNP+11].
@@
Moreover, $C_{c},C_{qa},C_{qc},C_{ns}$ are all closed, while $C_{q}$ and $C_{qa}$ are in general not.
3.3. Values of games
@@
4. Compiled nonlocal games
In this section we will review the construction from [KLVY23] that allows compiling any multi-player nonlocal game into a single-prover interactive protocol, along with the required cryptography. We will then prove a technical result that will be key to our later analysis. It states that as the security parameter tends to infinity, the average state of the prover after the first round of protocol becomes independent of the first round’s challenge (question) in a precise sense (Proposition 4.6).
4.1. Quantum homomorphic encryption
@@
It follows from [KLVY23, NZ23] that the quantum fully homomorphic encryption schemes of [Mah20, Bra18] can be used to define QHE schemes in the sense of the above definition (note that we require correctness only for a single $C\in\mathcal{C}$ at a time, as the security parameter tends to infinity).
@@
The security property demands that no adversary described by a QPT algorithm [^4] QPT algorithms as defined in Section 2 are a uniform notion. A stronger requirement is security against non-uniform QPT quantum adversaries, in which case one can also hope to get stronger conclusions. This is indeed the case and we return to this point in Remark 4.5 below. can distinguish between the encryption of any two fixed messages, with non-negligible probability, even when given access to an encryption oracle. This in fact implies a (seemingly stronger) security property, called parallel repeated IND-CPA security, where the adversary can choose the two messages and also receives a polynomial number of ciphertexts [NZ23].
4.2. The KLVY compiler
We now describe the compiler of [KLVY23]. It takes as its input a nonlocal game (Definition 3.1) and a QHE scheme (Definition 4.1). We assume from here onwards that the question and answer sets of the game are encoded as bitstrings of some fixed length.
Definition 4.2 ([KLVY23]).
@@ 4.2. The KLVY compiler
We only described the compiled version of a two-player nonlocal game, which is the focus of the present work, but the compiler generalizes straightforwardly to any game with $k$ players (in which case $2k$ rounds of communication are required) [KLVY23].
@@
(4.1) $\omega_{\lambda}(\mathcal{G}_{\textnormal{comp}},S)=\omega(\mathcal{G},p_{\lambda})=\sum_{x\in\mathcal{I}_{A},y\in\mathcal{I}_{B}}\sum_{a\in\mathcal{O}_{A},b\in\mathcal{O}_{B}}\mu(x,y)V(a,b|x,y)p_{\lambda}(a,b|x,y),$
@@
Just like in the case of the nonlocal games, there are different scenarios, depending on whether we consider classical or quantum provers. Here we focus on the quantum scenario, since [KLVY23] already proved that no classical efficient prover can exceed the classical value of the nonlocal game. The following definition describes the behavior of an efficient quantum prover in the compiled setting, analogously to Definition 3.4 in the nonlocal setting.
@@
subject to QPT assumptions that are less straightforward to state than above. The relation is immediate: we take $\lvert\psi_{\lambda}\rangle$ to be the all-zeros states on a suitable multi-qubit Hilbert space $\mathcal{H}_{\lambda}$ (but it can be any state that can be prepared by a QPT algorithm), the projective measurements $P^{\lambda}_{\xi\alpha}$ correspond to the first part of Definition 4.3, the unitaries $U^{\lambda}_{\xi\alpha}$ can be taken as the identity, [^5] This is without loss of generality: the unitaries $U^{\lambda}_{\xi\alpha}$ can always be absorbed into the second POVM, as $\lvert\alpha\rangle$ is part of the post-measurement state and we can also keep a copy of $\lvert\xi\rangle$ in it. and the operators $B^{\lambda}_{yb}$ correspond to the second part of Definition 4.3, that is,
(4.2) $B^{\lambda}_{yb}=(\langle y\rvert\otimes\mathbbm{1})W_{\lambda}^{*}(\lvert b\rangle\langle\mathopen{}b\rvert\otimes\mathbbm{1})W_{\lambda}(\lvert y\rangle\otimes\mathbbm{1}).$
@@
(4.3) $\sigma_{xa}^{\lambda}=\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{\xi\leftarrow\mathsf{Enc}(\mathsf{sk},x)}\sum_{\alpha\in\mathsf{CT}}\Pr(a\leftarrow\mathsf{Dec}(\mathsf{sk},\alpha))\,A_{\xi\alpha}^{\lambda}\lvert\psi^{\lambda}\rangle\langle\mathopen{}\psi^{\lambda}\rvert(A_{\xi\alpha}^{\lambda})^{*},$
@@
(4.4) $p_{\lambda}(a,b|x,y)$ $=\operatorname{tr}(\sigma^{\lambda}_{xa}B_{yb}^{\lambda}).$
@@
Then we associate to of a compiled game $\mathcal{G}$ the following minimal and maximal quantum values:
@@
Any strategy for the nonlocal game can be converted into a prover for the compiled game by using the homomorphic evaluation functionality of the encryption (assuming it supports evaluating the necessary quantum circuits). Thus, for every quantum strategy $S$ for the nonlocal game $\mathcal{G}$, there exists a QPT strategy $S_{\textnormal{comp}}=\{S_{\textnormal{comp}}^{\lambda}\}$ and a negligible function $\eta$ such that $\omega_{\lambda}(\mathcal{G}_{\textnormal{comp}},S_{\textnormal{comp}})\geq\omega(\mathcal{G},S)-\eta(\lambda)$ for all $\lambda\in\mathbb{N}$. This is one half of the main result [KLVY23, Thm. 3.2], and it implies that in particular $\omega_{q,\min}(\mathcal{G}_{\textnormal{comp}})\geq\omega_{q}(\mathcal{G})$. The other half of their theorem states that efficient classical provers cannot exceed the classical value $\omega_{c}(\mathcal{G})$ by a non-negligible amount, as already mentioned earlier.
@@
In this section, we consider quantum strategies that are defined in terms of (uniform) QPT algorithms, in line with the prior works [NZ23, CMM+24] and to emphasize that all reductions that will be discussed in the following are uniform as well. In cryptography, one can also model adversaries by non-uniform QPT algorithms, as mentioned earlier, and one can similarly define non-uniform QPT strategies. It is easy to see that all our results hold verbatim for such strategies, provided the QHE scheme is secure against non-uniform QPT adversaries.
@@
4.3. Asymptotic security for any noncommutative polynomial

### v3

@@ 2.2 Algebras and representations
for all $A\in\mathscr{A}$. Moreover, $\lvert\nu_{\phi}\rangle$ is cyclic (meaning $\overline{\pi_{\phi}(\mathscr{A})\lvert\nu_{\phi}\rangle}=\mathcal{H}_{\phi}$) and thereby uniquely determined. We call $(\mathcal{H}_{\phi},\pi_{\phi},\lvert\nu_{\phi}\rangle)$ a GNS triple associated with $\phi$. For more information on $C^{*}$-algebras, we refer the reader to [5].
@@
2.3 Classical and quantum computing
@@
For quantum computations, we will use the quantum circuit model. Here, computations correspond to the application of quantum circuits, which are unitary operators that operate on the Hilbert space $\mathcal{H}=(\mathbb{C}^{2})^{\otimes k}$ of some number $k$ of qubits and are given by the composition of unitary gates that each act nontrivially only on (for definiteness) one or two qubits (taken from some fixed universal gate set). The size of a quantum circuit is the number of gates used in the computation (we assume all qubits are acted upon by at least one gate). The qubits are typically split into input qubits and auxiliary qubits, which are assumed to be initialized in the $\lvert 0\rangle$ state, unless stated otherwise. If a classical outcome is desired, a subset of the qubits is measured after the unitary circuit has been applied. A quantum polynomial-time (QPT) algorithm consists of a family of quantum circuits $\{C_{\lambda}\}_{\lambda\in\mathbb{N}}$ and a deterministic polynomial-time Turing machine that on input $1^{\lambda}$ outputs a description of $C_{\lambda}$. We can often interpret $\lambda$ as a problem size or as a security parameter.
@@
3 Nonlocal games and strategies
@@ 3.1 Nonlocal games
A (two-player) nonlocal game is a tuple $\mathcal{G}=(\mathcal{I}_{A},\mathcal{I}_{B},\mathcal{O}_{A},\mathcal{O}_{B},\mu,V)$ describing a scenario consisting of non-communicating players, Alice and Bob, interacting with a referee. In the game, the referee samples a pair of questions $(x,y)\in\mathcal{I}_{A}\times\mathcal{I}_{B}$ according to $\mu$, sending question $x$ to Alice and $y$ to Bob. Then, Alice (resp. Bob) returns answer $a$ (resp. $b$) to the referee, who computes the rule function $V$ on the question-answer pairs $(a,b,x,y)$ to determine if $V(a,b|x,y)=1$ they win, or $V(a,b|x,y)=0$ they lose. [^1] We use the notation $V(a,b|x,y)$ instead of $V(a,b,x,y)$ to emphasize that this represents the value of answers $a,b$ given questions $x,y$.
All the information about the game $\mathcal{G}$ is available to the players before the game. This allows them to decide on a strategy beforehand. However, once the game begins the players are not allowed to communicate. To the referee, the behavior of the players can be modeled by the probabilities $p(a,b|x,y)$ of answers $a,b$ given questions $x,y$ as determined by the strategy. The collection of numbers $\{p(a,b|x,y)\}_{a\in\mathcal{O}_{A},b\in\mathcal{O}_{B},x\in\mathcal{I}_{A},y\in\mathcal{I}_{B}}\in\mathbb{R}^{\mathcal{O}_{A}\times\mathcal{O}_{B}\times\mathcal{I}_{A}\times\mathcal{I}_{B}}$ is called a (bipartite) correlation. Thus, the probability of winning the game $\mathcal{G}$ under a strategy $S$, with correlations $p$, is given by
$\omega(\mathcal{G},S)=\omega(\mathcal{G},p)=\sum_{x\in\mathcal{I}_{A},y\in\mathcal{I}_{B}}\sum_{a\in\mathcal{O}_{A},b\in\mathcal{O}_{B}}\mu(x,y)V(a,b|x,y)p(a,b|x,y).$ (3.1)
@@ 3.2 Strategies and correlations
Quantum strategies of particular interest are the entangled strategies. This is because if the state in the quantum strategy is unentangled, then the resulting correlation is always classical. It is easy to see that $C_{c}\subseteq C_{q}$, and the inclusion is in general strict, as follows from the existence of nontrivial Bell inequalities [9, 29, 39].
The restriction to finite-dimensional quantum systems is not the only model. If one allows the Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ to be infinite-dimensional, one gets the set of spatial quantum correlations $C_{qs}$. Clearly, $C_{q}\subseteq C_{qs}$, and both sets are convex subsets $\mathbb{R}^{\mathcal{O}_{A}\times\mathcal{O}_{B}\times\mathcal{I}_{A}\times\mathcal{I}_{B}}$. It turns out that in general, the inclusion is strict [13] and neither set is closed [42, 16, 11, 32, 3]. Moreover, both sets have the same closure, denoted by $C_{qa}$ and named the set of quantum approximable correlations.
@@
A (quantum) commuting operator strategy [^2] These are also called quantum commuting strategies in parts of the literature. for a nonlocal game $\mathcal{G}$ consists of
@@
The set of commuting operator correlations $C_{qc}\subseteq\mathbb{R}^{\mathcal{O}_{A}\times\mathcal{O}_{B}\times\mathcal{I}_{A}\times\mathcal{I}_{B}}$ is always a closed convex subset [35]. Since every quantum strategy is a commuting operator strategy by properties of the tensor product, it follows that $C_{qa}\subseteq C_{qc}$. When the Hilbert space $\mathcal{H}$ is finite-dimensional then any commuting operator strategy can also be seen as an ordinary quantum strategy, but in general, this is not so. In fact, there exist commuting operator correlations which have no realization as a quantum strategy, and hence $C_{q}\subsetneq C_{qc}$ [41]. Whether the correlation sets $C_{qa}$ and $C_{qc}$ were the same became known as Tsirelson’s Problem. This was recently resolved in the celebrated work $\mathsf{MIP}^{*}=\mathsf{RE}$ [22] by the construction of a correlation in $C_{qc}$ with no realization in $C_{qa}$. In turn, this implied a negative resolution to Connes’ Embedding Problem, following [36, 18, 23].
@@
Moreover, $C_{c},C_{qa},C_{qc},C_{ns}$ are all closed, while $C_{q}$ and $C_{qs}$ are in general not.
3.3 Values of games
@@
4 Compiled nonlocal games
In this section we will review the construction from [24] that allows compiling any multi-player nonlocal game into a single-prover interactive protocol, along with the required cryptography. We will then prove a technical result that will be key to our later analysis. It states that as the security parameter tends to infinity, the average state of the prover after the first round of protocol becomes independent of the first round’s challenge (question) in a precise sense (Proposition 4.6).
4.1 Quantum homomorphic encryption
@@
It follows from [24, 33] that the quantum fully homomorphic encryption schemes of [26, 6] can be used to define QHE schemes in the sense of the above definition (note that we require correctness only for a single $C\in\mathcal{C}$ at a time, as the security parameter tends to infinity).
@@
The security property demands that no adversary described by a QPT algorithm [^4] QPT algorithms as defined in Section 2 are a uniform notion. A stronger requirement is security against non-uniform QPT quantum adversaries, in which case one can also hope to get stronger conclusions. This is indeed the case and we return to this point in Remark 4.5 below. can distinguish between the encryption of any two fixed messages, with non-negligible probability, even when given access to an encryption oracle. This in fact implies a (seemingly stronger) security property, called parallel repeated IND-CPA security, where the adversary can choose the two messages and also receives a polynomial number of ciphertexts [33].
4.2 The KLVY compiler
We now describe the compiler of [24]. It takes as its input a nonlocal game (Definition 3.1) and a QHE scheme (Definition 4.1). We assume from here onwards that the question and answer sets of the game are encoded as bitstrings of some fixed length.
Definition 4.2 ([24]).
@@ 4.2 The KLVY compiler
We only described the compiled version of a two-player nonlocal game, which is the focus of the present work, but the compiler generalizes straightforwardly to any game with $k$ players (in which case $2k$ rounds of communication are required) [24].
@@
$\omega_{\lambda}(\mathcal{G}_{\textnormal{comp}},S)=\omega(\mathcal{G},p_{\lambda})=\sum_{x\in\mathcal{I}_{A},y\in\mathcal{I}_{B}}\sum_{a\in\mathcal{O}_{A},b\in\mathcal{O}_{B}}\mu(x,y)V(a,b|x,y)p_{\lambda}(a,b|x,y),$ (4.1)
@@
Just like in the case of the nonlocal games, there are different scenarios, depending on whether we consider classical or quantum provers. Here we focus on the quantum scenario, since [24] already proved that no classical efficient prover can exceed the classical value of the nonlocal game. The following definition describes the behavior of an efficient quantum prover in the compiled setting, analogously to Definition 3.4 in the nonlocal setting.
@@
subject to QPT assumptions that are less straightforward to state than above. The relation is immediate: we take $\lvert\psi_{\lambda}\rangle$ to be the all-zeros state on a suitable multi-qubit Hilbert space $\mathcal{H}_{\lambda}$ (but it can be any state that can be prepared by a QPT algorithm), the projective measurements $P^{\lambda}_{\xi\alpha}$ correspond to the first part of Definition 4.3, the unitaries $U^{\lambda}_{\xi\alpha}$ can be taken as the identity, [^5] This is without loss of generality: the unitaries $U^{\lambda}_{\xi\alpha}$ can always be absorbed into the second POVM, as $\lvert\alpha\rangle$ is part of the post-measurement state and we can also keep a copy of $\lvert\xi\rangle$ in it. and the operators $B^{\lambda}_{yb}$ correspond to the second part of Definition 4.3, that is,
$B^{\lambda}_{yb}=(\langle y\rvert\otimes\mathbbm{1})W_{\lambda}^{*}(\lvert b\rangle\langle\mathopen{}b\rvert\otimes\mathbbm{1})W_{\lambda}(\lvert y\rangle\otimes\mathbbm{1}).$ (4.2)
@@
$\sigma_{xa}^{\lambda}=\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{\xi\leftarrow\mathsf{Enc}(\mathsf{sk},x)}\sum_{\alpha\in\mathsf{CT}}\Pr(a\leftarrow\mathsf{Dec}(\mathsf{sk},\alpha))\,A_{\xi\alpha}^{\lambda}\lvert\psi^{\lambda}\rangle\langle\mathopen{}\psi^{\lambda}\rvert(A_{\xi\alpha}^{\lambda})^{*},$ (4.3)
@@
$p_{\lambda}(a,b|x,y)$ $=\operatorname{tr}(\sigma^{\lambda}_{xa}B_{yb}^{\lambda}).$ (4.4)
@@
Then we associate to a compiled game $\mathcal{G}_{\textnormal{comp}}$ the following minimal and maximal quantum values:
@@
Any strategy for the nonlocal game can be converted into a prover for the compiled game by using the homomorphic evaluation functionality of the encryption (assuming it supports evaluating the necessary quantum circuits). Thus, for every quantum strategy $S$ for the nonlocal game $\mathcal{G}$, there exists a QPT strategy $S_{\textnormal{comp}}=\{S_{\textnormal{comp}}^{\lambda}\}$ and a negligible function $\eta$ such that $\omega_{\lambda}(\mathcal{G}_{\textnormal{comp}},S_{\textnormal{comp}})\geq\omega(\mathcal{G},S)-\eta(\lambda)$ for all $\lambda\in\mathbb{N}$. This is one half of the main result [24, Thm. 3.2], and it implies that in particular $\omega_{q,\min}(\mathcal{G}_{\textnormal{comp}})\geq\omega_{q}(\mathcal{G})$. The other half of their theorem states that efficient classical provers cannot exceed the classical value $\omega_{c}(\mathcal{G})$ by a non-negligible amount, as already mentioned earlier.
@@
In this section, we consider quantum strategies that are defined in terms of (uniform) QPT algorithms, in line with the prior works [33, 15] and to emphasize that all reductions that will be discussed in the following are uniform as well. In cryptography, one can also model adversaries by non-uniform QPT algorithms, as mentioned earlier, and one can similarly define non-uniform QPT strategies. It is easy to see that all our results hold verbatim for such strategies, provided the QHE scheme is secure against non-uniform QPT adversaries.
@@
4.3 Asymptotic security for any noncommutative polynomial